OpenStudy (factor):

Resistors R1,R2,R3 have the values 5ohms, 10 ohms 30 ohms respectively which have been connected to a battery of 12 V. Calculate (a) the current through each resistor (b the total current in the circuit c) the total circuit resistence

OpenStudy (anonymous):

I think c

OpenStudy (factor):

its not mcq ur supposed to solce a b and c

OpenStudy (anonymous):

Ohhh okay so do you know Ohm's law?

OpenStudy (anonymous):

I'm assuming the circuit is in parallel. Let's find total resistance. if it's series, it's just a matter of adding em up, which gives \[5+10+30=45\] Let's find the TOTAL current first, which is according to \[V=IR\] So \[12V=I times 45\] \[I = 0.27A\] Note that since there is only one path for the current to travel, the current through each of the resistors is the same. Therefore your answers to part A and B both are 0.27A.

OpenStudy (anonymous):

Sorry, I assumed the circuit was in series*. Not parallel.

OpenStudy (anonymous):

The answer to C, total resistance, is as calculated above, 45 ohms.

OpenStudy (anonymous):

We need to know how the resistors are connected before we can answer the question. Was there a diagram with the question ?

OpenStudy (factor):

umm no

OpenStudy (radar):

From the context and the questions, I am believing they are connected in parallel. "Why" you may ask? First it implies that the resistors (all three) are connected to a battery. For each resistor to be connected to a battery it would be parallel because if in series one resistor would not have a connection to the battery.The total resistance R=E/I = 12/4 = 3 amps.